### v3

@@ 5.1 Implementation of Algorithm 1
The core of Algorithm 1 is the inexact computation of subproblem (14), the minimization of the sum of the $q$-order regularized quadratic approximation of $f$ at $x^{k}$ and the nonsmooth convex $g$. According to the discussion in Remark 3.4, we choose ZeroFPR as a solver to subproblems. ZeroFPR is a line-search algorithm proposed in [44] for the composite problem of form (1) with a general nonsmooth $g$ by seeking a critical point of the forward-backward envelope of $F$ with the limited-memory BFGS (lbfgs). When calling ZeroFPR to solve (14), we adopt its default setting except that the inexactness criterion (15) is used as the stopping condition and the maximum number of iterations is set as $10^{3}$. For the subsequent tests, the parameters of Algorithm 1 are chosen as follows
@@ 5.2 Validation of local convergence rate
We validate the local superlinear convergence rate and global complexity bound results of Algorithm 1. Figure 1 (a) plots the KKT residual curves yielded by Algorithm 1 with different $q\in(2,3]$ for solving the $\ell_{1}$-norm regularized logistic regression with dataset “leu”. For this class of convex composite problems, Section 5.3.1 contains its detailed description, and the local Lipschitzian error bound was shown to hold at any $\overline{x}\in\mathcal{S}^{*}=\mathcal{X}^{*}$ in [51]. We see that the KKT residual sequence returned by Algorithm 1 with a larger $q\in(2,3]$ indeed has a better $Q$-superlinear rate and global complexity bound, which coincides with the theoretical results of Theorems 4.10 and 4.16, but the times listed in the legend indicates that Algorithm 1 with a larger $q\in(2,3]$ may not require less one. Figure 1 (b) plots the residual curves returned by Algorithm 1 with different $q\in(2,3]$ for solving a test example of portfolio decision with higher moments from Section 5.3.3. For this class of nonconvex composite problems, now it is unclear whether the local Hölderian error bound in Theorems 4.10 and 4.16 or even the one in (6) hold, because only a tough upper estimation is available for the regular normal cone to ${\rm gph}\,\partial g$, which prevents us achieving a criterion like Corollary 2.4 (iii). The curves of Figure 1 (b) show that the sequence produced by Algorithm 1 with $q\in(2,3]$ has no superlinear rate, but Algorithm 1 with a larger $q$ still has a better global complexity bound and requires less running time according to those listed in the legend.
@@ 5.3.2 $\ell_{1}$-norm regularized Student’s $t$-regressions
This class of problem takes the form of (1) with $X=\mathbb{R}^{n},\,f(\cdot)\!=\psi(A\,\cdot\!-\!b)$ and $g(\cdot)\!=\lambda\|\cdot\|_{1}$, where $\psi(u)\!:=\!\sum_{i=1}^{m}\log(1+{u_{i}^{2}}/{\nu})\ (\nu>0)$ for $u\in\mathbb{R}^{m}$. Such a nonconvex loss function $f$ was introduced in [2] to deal with the data contaminated by heavy-tailed Student-$t$ noise. The test examples are randomly generated in the same way as in [32]. Specifically, we generate a true sparse signal $x^{\rm true}$ of length $n=512^{2}$ with $s=\lfloor\frac{n}{40}\rfloor$ nonzero entries whose indices are chosen randomly, and then compute each nonzero component by the formula $x^{\rm true}_{i}=\eta_{1}(i)10^{d\eta_{2}(i)/20}$, where $\eta_{1}(i)\in\{1,-1\}$ is a random sign and $\eta_{2}(i)$ is uniformly distributed in $[0,1]$. The signal has a dynamic range of $d$ dB with $d\in\{20,40,60,80\}$. The matrix $A\in\mathbb{R}^{m\times n}$ takes $m={n}/{8}$ random cosine measurements, i.e., $Ax=(\verb"dct"(x))_{J}$, where $dct"$ is the discrete cosine transform and $J\subseteq\{1,2,\ldots,n\}$ with $|J|=m$ is an index set chosen randomly. The vector $b$ is obtained by adding Student’s $t$-noise with degree of freedom $4$ and rescaled by $0.1$ to $Ax^{\rm true}$. We choose $\nu=0.25$ and $\lambda=c_{\lambda}\|\nabla\!f(0)\|_{\infty}$ with $c_{\lambda}>0$ for numerical tests.
We run three solvers from the starting point $x^{0}\!=A^{\top}b$ with the stopping accuracy $\epsilon=10^{-5}$ for $10$ independent trials, i.e., ten groups of data $(A,b)$ generated randomly. Table 3 reports the average objective values, KKT residuals, and running time over $10$ independent trials. We see that the three solvers produce the same objective values for all test examples, and our CR method and IRPNM return comparable KKT residuals for all test examples, which are superior to those returned by ZeroFPR. Compared with IRPNM, our CR method requires fewer iterations and needs less running time for $d=80$ but a little more time for $d=20,40$ and $60$. Notice that the test examples with $d=80$ are more difficult than those with $d=20,40$ and $60$. Moreover, after checking, we find that our CR method returns second-order stationary points for all test examples except for the case where $d=20,c_{\lambda}=0.1$ and the returned solutions $x^{*}$ for all test examples satisfy condition (b) of Corollary 2.4 (iii). Among others, the ${\rm H}_{\min}$ column in Table 3 reports the smallest one among the minimum eigenvalues of $[\nabla^{2}\!f(x^{*})]_{JJ}$ for $10$ trials. This is compatible with the fact that the CR method has a better local superlinear convergence rate and global complexity bound in theory.
@@
$d$ $c_{\lambda}$ Iter Fval resi time ${\rm H}_{\min}$ Iter Fval resi time Iter Fval resi time
20 0.1 10.4 9.53e+3 8.24e-6 10.2 0.31 147.8 9.53e+3 9.15e-6 9.3 24.6 9.53e+3 8.80e-6 9.3
0.01 5.7 1.02e+3 8.71e-6 55.2 0.26 1014.0 1.02e+3 9.56e-6 62.8 10.8 1.02e+3 8.30e-6 25.0
40 0.1 10.6 2.38e+4 8.82e-6 37.0 0.27 440.8 2.38e+4 9.26e-6 27.6 17.5 2.38e+4 8.40e-6 24.6
0.01 6.3 2.40e+3 8.07e-6 133.8 0.25 3255.9 2.40e+3 9.63e-6 202.6 14.2 2.40e+3 7.24e-6 81.6
60 0.1 7.0 5.42e+4 8.89e-6 87.5 0.26 1544.5 5.42e+4 9.48e-6 95.9 23.2 5.42e+4 8.09e-6 55.1
0.01 9.2 5.42e+3 8.77e-6 255.1 0.25 6272.6 5.42e+3 9.62e-6 390.0 17.6 5.42e+3 8.21e-6 178.2
80 0.1 8.0 1.35e+5 8.51e-6 196.6 0.25 5277.2 1.35e+5 9.51e-6 329.9 109.6 1.35e+5 8.35e-6 233.6
0.01 18.2 1.35e+4 8.77e-6 635.2 0.25 10000 1.35e+4 2.99e-4 624.2 116.4 1.35e+4 6.92e-6 794.5
@@ 5.3.3 Portfolio decision with higher moments
We choose the weekly returns $R\in\mathbb{R}^{N\times T}$ of NASDAQ100 index during the period 26/08/2022-25/08/2023 as the test dataset, where $N$ is the total number of stocks and $T$ is the number of weeks during the selected period. The test data is obtained by the following two steps: (1) randomly select $n\leq N$ stocks from the dataset and compute the return rate of each selected stock via $\xi_{it}=100[\log(R_{i,t+1})-\log(R_{i,t})]$ for $t=1,2,\ldots,T\!-\!1$; (2) compute four sample moments of the selected $n$ stocks as in [39] by using the return rate and the weight vector $(\omega_{1},\omega_{2},\omega_{3},\omega_{4})=(0.29,0.21,0.4,0.1)$. Table 4 reports the numerical results of three solvers with the stopping accuracy $\epsilon=10^{-5}$ and the starting point $x^{0}=\frac{1}{n}(1,1,\ldots,1)^{\top}$. We see that the CR method is significantly superior to ZeroFPR in terms of running time, and it returns the comparable objective values and a little better KKT residuals for all test examples. For this class of problems, the computation of function and gradient values of $f$ is expensive. The time difference between CR method and ZeroFPR is attributed to the fact that Algorithm 1 needs fewer calls of function and gradient values of $f$. Our CR method has the comparable performance with IRPNM except that the former needs a little fewer iterations and returns a little better objective values. For this class of test problems, IRPNM with the default setting exhibits poor numerical performance, failing to return a solution with the desired accuracy within the maximum number of iterations. The results of IRPNM in Table 4 are obtained by modifying IRPNM with a tighter approximation as suggested in [31, Section 4.2] to construct the subproblem as well as choosing the initial penalty parameter of dALMSN again.
@@ Acknowledgments
The authors would like to thank the anonymous referees and the associated editor for their valuable suggestions and comments, which helped improve this paper.

### v4

@@ 5.1 Implementation of Algorithm 1
The core of Algorithm 1 is the inexact computation of subproblem (10), the minimization of the sum of the $q$-order regularized quadratic approximation of $f$ at $x^{k}$ and the nonsmooth convex $g$. Based on Remark 3.3 in the archive version of this work, we choose ZeroFPR as a solver to subproblems. ZeroFPR is a line-search algorithm proposed in [43] for the composite problem of form (1) with a general nonsmooth $g$ by seeking a critical point of the forward-backward envelope of $F$ with the limited-memory BFGS (lbfgs). When calling ZeroFPR to solve (10), we adopt its default setting except that the inexactness criterion (11) is used as the stopping condition and the maximum number of iterations is set as $10^{3}$. For the subsequent tests, the parameters of Algorithm 1 are chosen as
@@ 5.2 Validation of local convergence rate
We validate the local superlinear convergence rate and global complexity bound results of Algorithm 1. Figure 1 (a) plots the KKT residual curves yielded by Algorithm 1 with different $q\in(2,3]$ for solving the $\ell_{1}$-norm regularized logistic regression with dataset “leu”. For this class of convex composite problems, Section 5.3.1 contains its detailed description, and the local Lipschitzian error bound was shown to hold at any $\overline{x}\in\mathcal{S}^{*}=\mathcal{X}^{*}$ in [49]. We see that the KKT residual sequence returned by Algorithm 1 with a larger $q\in(2,3]$ indeed has a better $Q$-superlinear rate and global complexity bound, which coincides with the theoretical results of Theorems 4.10 and 4.16, but the times listed in the legend indicates that Algorithm 1 with a larger $q\in(2,3]$ may not require less one. Figure 1 (b) plots the residual curves returned by Algorithm 1 with different $q\in(2,3]$ for solving a test example of portfolio decision with higher moments from Section 5.3.3. For this class of nonconvex composite problems, now it is unclear whether the local Hölderian error bound in Theorems 4.10 and 4.16 or even the one in (6) hold. The curves of Figure 1 (b) show that the sequence produced by Algorithm 1 with $q\in(2,3]$ has no superlinear rate, but Algorithm 1 with a larger $q$ still has a better global complexity bound and requires less running time by those in the legend.
@@ 5.3.2 $\ell_{1}$-norm regularized Student’s $t$-regressions
This class of problem takes the form of (1) with $\mathbb{X}=\mathbb{R}^{n},\,f(\cdot)\!=\psi(A\,\cdot\!-\!b)$ and $g(\cdot)\!=\lambda\|\cdot\|_{1}$, where $\psi(u)\!:=\!\sum_{i=1}^{m}\log(1+{u_{i}^{2}}/{\nu})\ (\nu>0)$ for $u\in\mathbb{R}^{m}$. Such a nonconvex loss function $f$ was introduced in [2] to deal with the data contaminated by heavy-tailed Student-$t$ noise. The test examples are randomly generated in the same way as in [31]. Specifically, we generate a true sparse signal $x^{\rm true}$ of length $n=512^{2}$ with $s=\lfloor\frac{n}{40}\rfloor$ nonzero entries whose indices are chosen randomly, and then compute each nonzero component by the formula $x^{\rm true}_{i}=\eta_{1}(i)10^{d\eta_{2}(i)/20}$, where $\eta_{1}(i)\in\{1,-1\}$ is a random sign and $\eta_{2}(i)$ is uniformly distributed in $[0,1]$. The signal has a dynamic range of $d$ dB with $d\in\{20,40,60,80\}$. The matrix $A\in\mathbb{R}^{m\times n}$ takes $m={n}/{8}$ random cosine measurements, i.e., $Ax=(\verb"dct"(x))_{J}$, where $dct"$ is the discrete cosine transform and $J\subseteq\{1,2,\ldots,n\}$ with $|J|=m$ is an index set chosen randomly. The vector $b$ is obtained by adding Student’s $t$-noise with degree of freedom $4$ and rescaled by $0.1$ to $Ax^{\rm true}$. We choose $\nu=0.25$ and $\lambda=c_{\lambda}\|\nabla\!f(0)\|_{\infty}$ with $c_{\lambda}>0$ for numerical tests.
We run three solvers from the starting point $x^{0}\!=A^{\top}b$ with the stopping accuracy $\epsilon=10^{-5}$ for $10$ independent trials, i.e., ten groups of data $(A,b)$ generated randomly. Table 3 reports the average objective values, KKT residuals, and running time over $10$ independent trials. We see that the three solvers yield the same objective values for all test examples, and our CR method and IRPNM return comparable KKT residuals for all test examples, which are superior to those returned by ZeroFPR. Compared with IRPNM, our CR method requires fewer iterations and needs less running time for $d=80$ but a little more time for other $d$. Notice that the test examples with $d=80$ are more difficult than those with $d=20,40$ and $60$. Moreover, after checking, we find that our CR method returns second-order stationary points for all test examples except for the case where $d=20,c_{\lambda}=0.1$. Among others, the $\lambda_{\rm min}$ column in Table 3 reports the smallest one among the minimum eigenvalues of $[\nabla^{2}\!f(x^{*})]$ for $10$ trials. This agrees with the fact that the CR method has a better local superlinear convergence rate and global complexity bound in theory.
@@
$d$ $c_{\lambda}$ Iter Fval resi time $\lambda_{\rm min}$ Iter Fval resi time Iter Fval resi time
20 0.1 10.4 9.53e+3 8.24e-6 10.2 -1 147.8 9.53e+3 9.15e-6 9.3 24.6 9.53e+3 8.80e-6 9.3
0.01 5.7 1.02e+3 8.71e-6 55.2 0 1014.0 1.02e+3 9.56e-6 62.8 10.8 1.02e+3 8.30e-6 25.0
40 0.1 10.6 2.38e+4 8.82e-6 37.0 0 440.8 2.38e+4 9.26e-6 27.6 17.5 2.38e+4 8.40e-6 24.6
0.01 6.3 2.40e+3 8.07e-6 133.8 0 3255.9 2.40e+3 9.63e-6 202.6 14.2 2.40e+3 7.24e-6 81.6
60 0.1 7.0 5.42e+4 8.89e-6 87.5 0 1544.5 5.42e+4 9.48e-6 95.9 23.2 5.42e+4 8.09e-6 55.1
0.01 9.2 5.42e+3 8.77e-6 255.1 0 6272.6 5.42e+3 9.62e-6 390.0 17.6 5.42e+3 8.21e-6 178.2
80 0.1 8.0 1.35e+5 8.51e-6 196.6 0 5277.2 1.35e+5 9.51e-6 329.9 109.6 1.35e+5 8.35e-6 233.6
0.01 18.2 1.35e+4 8.77e-6 635.2 0 10000 1.35e+4 2.99e-4 624.2 116.4 1.35e+4 6.92e-6 794.5
@@ 5.3.3 Portfolio decision with higher moments
We choose the weekly returns $R\in\mathbb{R}^{N\times T}$ of NASDAQ100 index during the period 26/08/2022-25/08/2023 as the test dataset, where $N$ is the total number of stocks and $T$ is the number of weeks during the selected period. The test data is obtained by the following two steps: (1) randomly select $n\leq N$ stocks from the dataset and compute the return rate of each selected stock via $\xi_{it}=100[\log(R_{i,t+1})-\log(R_{i,t})]$ for $t=1,2,\ldots,T\!-\!1$; (2) compute four sample moments of the selected $n$ stocks as in [38] by using the return rate and the weight vector $(\omega_{1},\omega_{2},\omega_{3},\omega_{4})=(0.29,0.21,0.4,0.1)$. Table 4 reports the numerical results of three solvers with the stopping accuracy $\epsilon=10^{-5}$ and the starting point $x^{0}=\frac{1}{n}(1,1,\ldots,1)^{\top}$. We see that the CR method is significantly superior to ZeroFPR in terms of running time, and it returns the comparable objective values and a little better KKT residuals for all test examples. For this class of problems, the computation of function and gradient values of $f$ is expensive. The time difference between CR method and ZeroFPR is attributed to the fact that Algorithm 1 needs fewer calls of function and gradient values of $f$. Our CR method has the comparable performance with IRPNM except that the former needs a little fewer iterations and returns a little better objective values. For this class of test problems, IRPNM with the default setting exhibits poor numerical performance, failing to return a solution with the desired accuracy within the maximum number of iterations. The results of IRPNM in Table 4 are obtained by modifying IRPNM with a tighter approximation as suggested in [30, Section 4.2] to construct the subproblem as well as choosing the initial penalty parameter of dALMSN again.
@@ Acknowledgments
The authors would like to thank the anonymous referees and the associated editor for their valuable suggestions and comments.